### v1

# Multimodal-enhanced Federated Recommendation: A Group-wise Fusion Approach

Abstract.
Federated Recommendation (FR) is a new learning paradigm to tackle the learn-to-rank problem in a privacy-preservation manner. How to integrate multi-modality features into federated recommendation is still an open challenge in terms of efficiency, distribution heterogeneity, and fine-grained alignment. To address these challenges, we propose a novel multimodal fusion mechanism in federated recommendation settings (GFMFR). Specifically, it offloads multimodal representation learning to the server, which stores item content and employs a high-capacity encoder to generate expressive representations, alleviating client-side overhead. Moreover, a group-aware item representation fusion approach enables fine-grained knowledge sharing among similar users while retaining individual preferences. The proposed fusion loss could be simply plugged into any existing federated recommender systems empowering their capability by adding multi-modality features. Extensive experiments on five public benchmark datasets demonstrate that GFMFR consistently outperforms state-of-the-art multimodal FR baselines.
Federated Learning, Multimodal Recommender Systems
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: $X$; 978-1-4503-XXXX-X/2018/06
1. Introduction
Driven by rising privacy concerns, Federated Recommendation (FR) has emerged as a prominent paradigm that enables collaborative model training without accessing raw user data, drawing increasing attention (Javeed et al., 2023; Sun et al., 2024; Wang et al., 2024). Despite this momentum, most FR methods rely heavily on structured item identifiers, such as unique IDs or tabular attributes, to model preferences. This narrow focus limits the ability to capture rich semantics from real-world items, which often include multimodal content like text, images, and audio(Liu et al., 2024a; Malitesta et al., 2025; Yang et al., 2025b). This substantially limits FR’s applicability, especially in domains where content signals play a central role in capturing user preferences(Liu et al., 2024b; Malitesta et al., 2024; Liu et al., 2023).
The effective integration of multimodal information into FR remains underexplored, primarily due to several inherent challenges. First, multimodal item attributes often have large raw data sizes and require complex processing for semantic extraction. This imposes substantial storage and computation burdens on resource-constrained client devices. Second, heterogeneous data distributions and varying modality usage across clients make it difficult to learn representations that generalize well while preserving personalization. Third, aligning content features with recommendation objectives requires more than basic feature fusion, demanding explicit guidance to avoid redundancy and noise. These challenges highlight the need for a principled framework capable of systematically integrating multimodal knowledge into FR models.
To this end, we introduce a conceptual framework named GFMFR (Group-wise Fusion for Multimodal-enhanced Federated Recommendation). The framework proposes a reallocation of learning responsibilities: Multimodal representation learning takes place centrally on the server side, where comprehensive item content and sufficient computational capacity are accessible. Clients then operate on compact, distilled representations that encapsulate both semantic content and user preference relevance. This design reduces client-side burdens and aligns well with commonly adopted system architectures in real-world deployments.
In addition to improving system efficiency, GFMFR incorporates a group-aware multimodal aggregation mechanism, which enables shared representation learning among users with similar preferences. This facilitates cross-client knowledge transfer while maintaining personalized modeling capabilities. Moreover, a preference-guided distillation strategy is introduced to ensure that multimodal signals directly contribute to the learning of recommendation-relevant features. Extensive experiments demonstrate that our method achieves state-of-the-art performance on benchmark datasets. Our main contributions are summarized as follows,
  • We propose GFMFR, a novel conceptual framework that systematically integrates multimodal content into federated recommendation, providing a flexible and effective approach to leverage diverse item modalities.
  • We design a group-aware multimodal aggregation mechanism to enable fine-grained integration of multimodal representations among users with similar preferences. Additionally, we incorporate a preference-guided distillation loss to align multimodal features with recommendation objectives, balancing knowledge sharing and personalization.
  • Extensive experiments demonstrate the superior performance of GFMFR compared to competitive baselines. In addition, comprehensive analyses further validate its compatibility, effectiveness and practical applicability.
2. Related Work
2.1. Multimodal Recommender System
With the proliferation of multimedia platforms and diverse content (e.g., short videos, images, news), incorporating multimodal information into recommender systems has become crucial for capturing user preferences and item characteristics. Such integration not only enriches representations but also alleviates data sparsity issues (Liu et al., 2024a; Zhang et al., 2024a; Zhao et al., 2025), making multimodal recommendation a prominent research direction (Ye et al., 2025; Zhuang et al., 2025; Zhou et al., 2025). The typical pipeline involves three stages: raw feature representation (Javed et al., 2021; Devlin et al., 2019), feature fusion (Wang et al., 2021; Tao et al., 2022; Yang et al., 2025a), and recommendation prediction (Koren et al., 2009; Meng et al., 2025). Raw feature extraction aims to obtain informative modality-specific representations (Chen et al., 2022; He et al., 2016); fusion methods (Zhou et al., 2023; Ma et al., 2022) integrate them into a unified space; and prediction models (He and McAuley, 2016; Xu et al., 2025) generate personalized recommendations. Despite their effectiveness, existing centralized approaches raise privacy concerns, limiting their applicability. To address this, we propose a multimodal recommendation framework under federated learning, ensuring privacy protection while retaining recommendation quality.
2.2. Federated Recommender System
Federated recommendation (FR) offers personalized services while safeguarding user privacy, making it well-suited for privacy-sensitive scenarios (Alamgir et al., 2022; Chronis et al., 2024; Harasic et al., 2024; Li and Long, 2024; Zhang et al., 2025b). Existing FR research spans two major perspectives: recommender systems and federated learning. From the recommender perspective, studies have explored diverse model architectures, including MF (Ammad-Ud-Din et al., 2019; Chai et al., 2020), MLP (Muhammad et al., 2020; Perifanis and Efraimidis, 2022), GNN (Wu et al., 2022; Zhang et al., 2024b), and Transformer (Wei et al., 2023; Zhang et al., 2025a), as well as application scenarios such as cross-domain (Meihan et al., 2022; Zhang et al., 2024d), cold-start (Zhang et al., 2024c; Li et al., 2025), and news recommendation (Qi et al., 2020; Yi et al., 2021). From the federated learning perspective, efforts have addressed core challenges in security (Liu et al., 2022; Perifanis et al., 2023), robustness (Zhang et al., 2022; Ali et al., 2024), and efficiency (Zhang et al., 2023b; Nguyen et al., 2024). However, most FR methods still rely on ID-based item representations, overlooking rich multimodal content (e.g., text, images), which can offer valuable semantic cues.
Research on multimodal FR remains limited. Existing methods such as AMMFRS (Feng et al., 2024) and FedMR (Li et al., 2024c) adopt client-side fusion strategies, integrating multimodal features locally: AMMFRS via attention-based weighting and FedMR via interaction-aware fusion modules. However, these approaches assume clients can store and process large-scale multimodal data, which is impractical for resource-constrained devices. FedMMR (Li et al., 2024a) shifts fusion to the server and aligns modalities using a contrastive loss, but focuses primarily on cross-modal alignment while overlooking inter-modal semantic dependencies. To address these limitations, we propose a novel server-side multimodal fusion framework that offloads both storage and representation learning. By further incorporating inter-user correlations, our approach integrates multimodal signals in a fine-grained manner, improving recommendation quality while maintaining efficiency.
3. Problem Formulation
In this section, we present a formalization of our proposed conceptual framework for integrating multimodal content into federated recommender systems. We begin with the general formulation of FR, followed by a clear clarification of the problem and a detailed explanation of the corresponding optimization objective.
General FR framework. Let $\mathcal{U}$ and $\mathcal{I}$ denote the user set and item set in the recommender system, and $N=|\mathcal{U}|$ and $M=|\mathcal{I}|$ are the total number of users and items respectively. For a recommendation model $\mathcal{F}$ parameterized by $\theta$, it is generally composed of three modules, i.e., user embedding $E_{\mathcal{U}}$, item embedding $E_{\mathcal{I}}$ and prediction function $P$. Given user $u$ and item $i$, the recommendation model makes rating predictions as $\hat{\mathcal{R}}_{ui}=\mathcal{F}(u,i|\theta)$. In the federated recommender system, each user $u$ acts as a client who holds the personal data $\mathcal{R}_{u}$, including all the historically rated items. To implement an FR model, the system alternates between the clients and server and performs the following steps:
(1) Client-side model training: The client $u$ trains the locally maintained recommendation model $\mathcal{F}_{\theta}$ with personal data $\mathcal{R}_{u}$. (2) Server-side model aggregation: A server collects the updated models from clients and performs a global aggregation operation to achieve a unified model. (3) Global model synchronization: The unified model would be sent to the clients to improve their models’ training in the next round. Formally, the optimization objective of FR can be formulated as follows,
(1) $\min_{\{E_{\mathcal{U}}^{u}\}_{u=1}^{N},E_{\mathcal{I}},P}\sum_{u=1}^{N}\alpha_{u}\mathcal{L}_{rec}(\theta;\mathcal{R}_{u})$
where $\mathcal{L}_{rec}$ is local recommendation model loss of the $u$-th client, and $\alpha_{u}$ is the weight of the $u$-th client during global aggregation. For example, $\alpha_{u}$ is usually proportional to the client’s personal data volume and the total system data volume (McMahan et al., 2017). In existing FR research, the user embedding module $E_{\mathcal{U}}$ is typically treated as a private component and retained locally on each client, while the item embedding $E_{\mathcal{I}}$ and prediction function $P$ are shared with the server for global collaborative training.
Figure 1. Overview of the proposed GFMFR. The left part illustrates a general FR framework. Given $k$ modalities of raw input data $\{X_{m}\}_{m=1}^{k}$, GFMFR builds upon this by introducing an additional fusion loss $\mathcal{L}_{fus}$, which integrates the ID-based feature $E_{\mathcal{I}}$ learned from the recommendation model $\mathcal{F}_{\theta}$ with multimodal feature $E_{\mathcal{M}}$ learned by a multimodal representation model $\mathcal{H}_{\psi}$, thereby enabling multimodal information integration into the federated recommendation system.
The proposed multimodal-enhanced FR framework. Given raw multimodal item data $\{X_{m}\}_{m=1}^{k}$, where $k$ is the number of item modalities. Let $\mathcal{H}_{\psi}$ denotes the multimodal representation learning model, and $E_{\mathcal{M}}=\mathcal{H}_{\psi}(\{X_{m}\}_{m=1}^{k})$ denotes the learned multimodal representation. The multimodal-enhanced FR framework aims to integrate the multimodal representation $E_{\mathcal{M}}$ into the item embedding $E_{\mathcal{I}}$ through a dedicated fusion loss $L_{fus}$, thereby improving the overall effectiveness of FR learning. Formally, we formulate the optimization objective of the multimodal-enhanced FR framework as,
(2) $\min_{\{E_{\mathcal{U}}^{u}\}_{u=1}^{N},E_{\mathcal{I}},P}\sum_{u=1}^{N}\alpha_{u}(\mathcal{L}_{rec}(\theta;\mathcal{R}_{u})+\lambda\cdot\mathcal{L}_{fus}(E_{\mathcal{I}},E_{\mathcal{M}}))$
where $\lambda$ is the coefficient of fusion loss. The proposed multimodal-enhanced FR framework is inherently general and compatible, allowing for straightforward integration into existing FR models with minimal architectural modifications. The framework overview is illustrated in Figure 1.
4. Group-wise Fusion for Multimodal-enhanced FR (GFMFR)
In this section, we introduce the proposed Group-wise Fusion for Multimodal-enhanced FR (GFMFR). The framework delegates multimodal representation learning to the server, aligning with practical deployments where intensive computations are handled centrally. To enhance cross-client knowledge sharing, we design a group-aware aggregation mechanism enabling shared representation among users with similar preferences. A preference-guided distillation strategy further maps multimodal features into the preference space and distills them into local models, enabling effective integration of multimodal signals into FR. The following subsections detail the framework’s components.
4.1. Multimodal Representation Initialization
In multimodal recommendation, item profiles comprise diverse modalities such as text, images, and audio. To leverage these signals, raw inputs from each modality are encoded into model-compatible representations via modality-specific encoders. The server employs a powerful multimodal representation model to produce initial embeddings $\mathcal{E}_{m}\in\mathbb{R}^{M\times d_{1}}$ for $m=1,\ldots,k$, where $d_{1}$ denotes the embedding dimension. Leveraging abundant server resources, large-scale pre-trained models serve as backbones to generate high-quality initial representations.
4.2. Group-Aware Multimodal Aggregation
In FR systems, heterogeneous data distributions and modality usage across clients cause inconsistent multimodal representations, impeding knowledge sharing. Global representation enforcement improves consistency but compromises personalization. To address this, we propose a group-aware multimodal aggregation that derives shared representations for preference-aligned user groups, enabling efficient knowledge transfer while preserving client-specific modeling.
Cluster clients into groups. Clients are clustered by their prediction functions, which capture decision patterns over items and group users with similar preferences. This method preserves privacy by eliminating the need to access raw user data. Specifically, the server aggregates clients’ prediction functions $\{P_{u}\}_{u=1}^{N}$ and clusters them based on their parameters, formally expressed as,
(3) $G_{1},G_{2},...,G_{g}=\text{ClusterAlgorithm}(\{P_{u}\}_{u=1}^{N})$
where $g$ denotes the total number of user groups, and common choices for the clustering algorithm include K-Means and hierarchical clustering.
Learn group-level multimodal features. Given initial representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ and user groups $\{G_{1},...,G_{g}\}$, a multimodal aggregation module learns shared features within each group. This module comprises a learnable group embedding $E_{G}\in\mathbb{R}^{g\times d_{2}}$ assigning unique IDs to groups, and an aggregation function $S_{\phi}$. For each group $l=1,\dots,g$, the aggregation is formulated as,
(4) $Q_{l}=S_{\phi}(\{\mathcal{E}_{m}\}_{m=1}^{k},E_{G}(l))$
where $E_{G}(l)$ is the embedding for group $l$, and $Q_{l}$ denotes the aggregated multimodal features. To optimize this module, we use group-level item embeddings as supervision. Specifically, for group $l$, the item embedding $C_{l}$ is obtained by aggregating local embeddings from members, e.g., $C_{l}=\text{AvgPooling}(\{E_{\mathcal{I}}^{u}\}_{u\in G_{l}})$. The parameters of $S_{\phi}$ are then updated by minimizing,
(5) $\min_{E_{G},\phi}\frac{1}{g}\sum_{l=1}^{g}\mathcal{L}_{agg}(E_{G},\phi;C_{l})$
where $\mathcal{L}_{agg}$ is the aggregation loss, e.g., the mean squared error between $Q_{l}$ and $C_{l}$. This optimization aligns aggregated group-level multimodal features with group-specific item embeddings, enhancing cross-modal consistency and yielding representations better suited for downstream recommendation.
4.3. Preference-Guided Distillation
To effectively incorporate multimodal features, we propose a preference-guided distillation strategy that projects group-level multimodal features into the preference space, providing auxiliary supervision for client-side training to enable goal-aligned knowledge transfer and enhanced personalization.
Map group-level multimodal features into preference space. Given the group-level multimodal features $\{Q_{l}\}_{l=1}^{g}$, we map them into the preference space via aggregated group prediction functions. Specifically, we obtain the group prediction function by averaging client functions within each group, $\overline{P}_{l}=\text{AvgPooling}(\{P_{u}\}_{u\in G_{l}})$, and apply $\overline{P}_{l}$ to $Q_{l}$ to generate the group’s preference representation,
(6) $E_{\mathcal{M}\_l}=\overline{P}_{l}(Q_{l})$
Once the group’s preference representations $\{E_{\mathcal{M}\_l}\}_{l=1}^{g}$ are obtained, the server would distribute them to clients for improving the local model’s training.
Distill group preference to client recommendation models. Each client trains its recommendation model locally on private data. Upon receiving the group-level preference representation from the server, it incorporates this as auxiliary supervision to refine the local model. This distillation transfers multimodal and group-aligned knowledge to enhance semantic capture while preserving personalization. Formally, the distillation objective for client $u$ in group $l$ is,
(7) $\min_{\theta_{u}}\mathcal{L}_{dis}(\theta_{u};E_{\mathcal{M}\_l})$
where $\theta_{u}$ denote the $u$-th client’s model parameters, and $\mathcal{L}_{dis}$ is the distillation loss function, e.g., the KL divergency between model’s prediction $\hat{\mathcal{R}}_{u}$ and group preference $E_{\mathcal{M}\_l}$.
4.4. Optimization Objective and Algorithm
Optimization objective. GFMFR is a multimodal FR system that distills group-aware preference signals from multimodal item semantics to guide personalized client model training. Based on the above procedure, we reformulate the overall optimization objective in Eq. (2) as follows,
(8) $\min_{\theta}\sum_{u=1}^{N}\alpha_{u}\left(\mathcal{L}_{rec}(\theta;\mathcal{R}_{u})+\lambda\cdot\mathcal{L}_{dis}(\theta;E_{\mathcal{M}\_l})\right)$
$\text{s.t.}\quad(E_{G}^{*},\phi^{*})=\arg\min_{E_{G},\phi}\frac{1}{g}\sum_{l=1}^{g}\mathcal{L}_{agg}(E_{G},\phi;C_{l}),$
$\quad Q_{l}=S_{\phi^{*}}(\{\mathcal{E}_{m}\}_{m=1}^{k},E_{G}^{*}(l)),\quad E_{\mathcal{M}\_l}=\overline{P}_{l}(Q_{l})$
where $\lambda$ is the coefficient of distillation loss. $(E_{G}^{*},\phi^{*})$ are the optimized parameters of the multimodal aggregation module, used to generate group-level multimodal representations $Q_{l}$, which are subsequently mapped to preference representations $E_{\mathcal{M}\_l}$. Here, $l$ denotes the group index assigned to the $u$-th client, which is dynamically updated during the federated optimization process. The detailed GFMFR procedure is illustrated in Appendix.
Algorithm. To solve the above optimization objective, we conduct an iterative algorithm between the server and clients to learn the multimodal FR system. As illustrated in Appendix, before the federated optimization begins, the server first performs a unified initialization of the globally shared parameters, including the item embedding $E_{\mathcal{I}}$ and the prediction function $P$. In parallel, high-quality initial multimodal representations $\{\mathcal{E}_{m}\}_{m=1}^{k}$ are obtained using a powerful multimodal representation learning model. In each communication round $\bm{t}$, selected clients $u\in B^{(t)}$ initialize local models with the latest shared parameters $E_{\mathcal{I}}^{(t)}$ and $P^{(t)}$, perform local updates using personal data $\mathcal{R}_{u}$, and upload the updated shared components to the server. Upon receiving the updated shared parameters $\{E_{\mathcal{I}}^{u}\}_{u\in B^{(t)}}$ and $\{P_{u}\}_{u\in B^{(t)}}$, the server first performs global aggregation to obtain the new shared parameters $E_{\mathcal{I}}^{(t+1)}$ and $P^{(t+1)}$. Concurrently, it clusters clients into groups $\{G_{1},G_{2},...,G_{g}\}^{(t)}$ based on the shared prediction functions $\{P_{u}\}_{u\in B^{(t)}}$, and trains a multimodal aggregation module with group-specific item embeddings $\{C_{l}^{(t)}\}_{l=1}^{g}$ to derive group-level multimodal representations $\{Q_{l}^{(t)}\}_{l=1}^{g}$. These representations are then transformed into group preference signals $\{E_{\mathcal{M}\_l}^{(t)}\}_{l=1}^{g}$ via the group-specific prediction function $\{\overline{P}_{l}^{(t)}\}_{l=1}^{g}$, which are subsequently distributed to corresponding clients $u\in B^{(t)}$ to refine their local recommendation models with the distillation loss $\mathcal{L}_{dis}$.
4.5. Discussion about Practical Applicability
Usage guideline. The proposed multimodal-enhanced FR framework emphasizes modularity and compatibility for seamless integration into existing FR pipelines. The multimodal representation learning is decomposed into three stages for practical deployment and customization. First, initial modality-specific representations can be obtained by directly using powerful pre-trained models on the server or trained from scratch, supporting flexible domain adaptation. Second, the group-aware multimodal aggregation module supports flexible grouping strategies (e.g., behavior, content, geography) and aggregation mechanisms (e.g., MLP (Rosenblatt, 1958), self-attention (Vaswani et al., 2017), mixture-of-experts (Jacobs et al., 1991)), with adjustable aggregation frequency balancing personalization and efficiency. Third, the preference-guided distillation can be replaced by alternative integration methods (e.g., adapter modules, attention fusion) to align multimodal knowledge with local objectives. These design choices ensure adaptability across diverse system constraints and scenarios.
Efficiency analysis. Multimodal item attributes, derived from raw data such as images, text, and audio, require high-capacity models for effective encoding. In GFMFR, storing and encoding these attributes is fully offloaded to the server, leveraging its superior computational and storage resources. This eliminates the burden on resource-limited clients from handling large multimodal data or complex models. The server further compresses rich multimodal semantics into low-dimensional preference representations (e.g., dimension 2 in implicit feedback tasks), which are then transmitted to clients to enhance local training. This approach preserves task-relevant semantics while significantly reducing communication overhead in federated optimization.
Security analysis. Unlike conventional ID-based representations, multimodal features are semantically rich but may inadvertently encode sensitive user attributes when learned locally, increasing the risk of privacy breaches via gradient inversion or attribute inference. In our framework, representation learning is entirely server-side, and clients receive only compact item preference vectors derived from multimodal features, which are less informative and more resilient to such attacks. This design mitigates privacy risks associated with multimodal signals in FR while remaining compatible with ID-based architectures, and amenable to integration with techniques like Differential Privacy (Choi et al., 2018) and Homomorphic Encryption (Acar et al., 2018). Further analysis is provided in the experiments.
5. Experiment
5.1. Experimental Setup
Dataset. We conduct experiments on five multimodal recommendation benchmark datasets from the Amazon collection(Hou et al., 2024), each representing a different item category (Beauty, Tools_and_Home, Toys_and_Games, Digital_Music, Office_Products), to evaluate the effectiveness of our proposed method. Detailed data descriptions are summarized in Appendix.
Baselines. Our proposed GFMFR is a general multimodal FR framework that can be seamlessly integrated with existing ID-based FR models. To verify its versatility, we select three advanced ID-based FR models as the backbones, including FedNCF (Perifanis and Efraimidis, 2022), PFedRec (Zhang et al., 2023a) and FedRAP (Li et al., 2024b). In addition, we include three existing multimodal FR models as baselines for comparison, including FedMMR, AMMFRS and FedMR. In addition, we extend a centralized multimodal recommendation model (Zhang et al., 2017) to the federated setting (named FedMLP) by employing an MLP-based fusion strategy on the client side, thereby enriching the comparison pool with a comprehensive coverage of commonly used multimodal fusion techniques. More details can be found in Appendix.
Evaluation protocol and implementation details. We evaluate the performance using two ranking-based metrics for top-k recommendation, i.e., Hit Rate (HR@k) and Normalized Discounted Cumulative Gain (NDCG@k). Specifically, we predict over the full item set and calculate the metrics based on the rank of the test item within the top-50 recommendation list. All experiments are conducted five times, with their average results consistently reported in the paper. More implementation details can be found in Appendix.
5.2. Overall Comparison
Overall performance. Table 1 reports HR@50 and NDCG@50 comparisons across five datasets. Key insights from the results are as follows:
(1) Integrating multimodal information into FR models can typically enhance system performance. By incorporating semantic cues, multimodal data enriches the system’s understanding of items. This deeper understanding enables more effective knowledge transfer between clients, which in turn boosts system performance. (2) Our method demonstrates superior performance compared to baselines. Existing multimodal FR methods focus on enhancing ID embeddings via specific fusion strategies, focusing primarily on feature enhancement. In contrast, our method leverages multimodal information to learn preference representations that are more aligned with the optimization objectives of recommendation models, leading to improved performance. (3) Our method showcases exceptional compatibility across various FR backbones. GFMFR serves as a general multimodal FR framework that can be seamlessly integrated with existing ID-based methods. Experimental results demonstrate that incorporating our model into three different FR backbones consistently yields the best performance across all datasets, highlighting its strong compatibility.
Table 1. Overall performance comparison of our proposed GFMFR and baselines integrated into three representative FR backbone models. “H@50” and “N@50” are abbreviations for HR@50 and NDCG@50. Bold indicates the best result, underline marks the second best result, and “*” denotes a statistically significant improvement over the best baseline (two-sided t-test, p ¡ 0.05).
Model Beauty Tools_and_Home Toys_and_Games Digital_Music Office_Products
H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50
FedNCF 0.1450 0.0407 0.1893 0.0483 0.2858 0.2436 0.2249 0.1822 0.0015 0.0004
w/ FedMMR 0.1331 0.0358 0.1198 0.0287 0.0354 0.0089 0.0200 0.0052 0.0130 0.0035
w/ AMMFRS 0.1549 0.0449 0.1947 0.0503 0.0733 0.0221 0.0417 0.0098 0.0145 0.0039
w/ FedMR 0.1601 0.0378 0.1777 0.0405 0.0565 0.0141 0.0267 0.0074 0.0142 0.0036
w/ FedMLP 0.1537 0.0364 0.1653 0.0444 0.0558 0.0120 0.0222 0.0053 0.0039 0.0011
[…]
PFedRec 0.1427 0.0320 0.1596 0.0443 0.1412 0.0626 0.1287 0.0728 0.0111 0.0027
w/ FedMMR 0.1383 0.0351 0.1212 0.0301 0.0401 0.0097 0.0228 0.0054 0.0118 0.0031
w/ AMMFRS 0.1594 0.0418 0.1791 0.0451 0.0829 0.0258 0.0455 0.0142 0.0139 0.0036
w/ FedMR 0.2022 0.0545 0.1756 0.0600 0.0524 0.0205 0.0612 0.0321 0.0164 0.0046
w/ FedMLP 0.1537 0.0361 0.1645 0.0417 0.1207 0.0516 0.0875 0.0361 0.0266 0.0074
[…]
FedRAP 0.1423 0.0363 0.1641 0.0560 0.0616 0.0226 0.0531 0.0273 0.0118 0.0033
w/ FedMMR 0.1331 0.0374 0.1291 0.0324 0.0397 0.0099 0.0224 0.0057 0.0103 0.0028
w/ AMMFRS 0.1522 0.0421 0.1384 0.0342 0.0414 0.0102 0.0182 0.0043 0.0160 0.0040
w/ FedMR 0.1568 0.0366 0.1896 0.0578 0.0480 0.0123 0.0280 0.0075 0.0112 0.0028
w/ FedMLP 0.1771 0.0479 0.1624 0.0562 0.0547 0.0225 0.0555 0.0298 0.0178 0.0051
[…]
Efficiency comparison. We further evaluate and compare the efficiency of our method and baselines from three aspects: (1) client-side storage overhead, including multimodal data representations (preferences) and recommendation model parameters; (2) communication overhead, covering the initial distribution of multimodal representations and the per-round transmission cost; and (3) per-epoch training time of the client-side recommendation model. As shown in Appendix, baselines AMMFRS, FedMR and FedMLP incur higher client storage overhead due to the need to store raw multimodal features and the integration of fusion modules within local models. Baseline FedMMR’s storage overhead primarily stems from its architecture by integrating an elaborate contrastive loss. These baselines also require longer client training time due to more complex local models. In terms of communication, baseline AMMFRS, FedMR and FedMLP transmit full initial multimodal representations to all clients before training, resulting in substantial communication overhead. In contrast, our method only exchanges low-dimensional group preference vectors (e.g., dimension 2 vs. 512), introducing minimal additional overhead. Overall, our method achieves higher efficiency and is better suited for real-world deployment constraints.
5.3. Ablation Study
This subsection examines the impact of the group-aware multimodal aggregation mechanism and the preference-guided distillation strategy on model performance. The analysis clarifies their individual contributions and demonstrates how each component enhances overall effectiveness. We conduct the evaluation by incorporating our method into PFedRec and reporting HR@50 results across all datasets. For results on NDCG@50, see Figure LABEL:fig:ablation_ndcg in Appendix.
Group-aware multimodal aggregation mechanism. We perform a step-wise ablation study with three key variants: replace preference-based clustering with random grouping to assess grouping quality (RandomGroup); treat all users as one group to examine the need for group differentiation (SingleGroup); and use group-specific aggregation modules instead of a shared one to evaluate the benefit of shared representation learning (MultipleAgg). As shown in Figure 2 (a), random grouping notably degrades performance, highlighting the importance of behavior-aware user partitioning. Treating all users as a single group also impairs results, as it overlooks individual preferences. Furthermore, using multiple aggregation modules reduces performance due to increased complexity and the absence of shared representation learning, compromising generalization and stability.
Preference-guided distillation strategy. The preference-enhanced distillation strategy leverages group-level preference representations to improve client training. To examine the effect of preference signal utilization, we compare four coefficient scheduling methods for the distillation loss: (1) match the scale of recommendation loss (ScaleAlign), (2) smooth with the previous round’s coefficient (SmoothCoef), (3) progressively increase the coefficient (ProgressiveScale), and (4) our strategy combines (2) and (3) (Ours). As shown in Figure 2 (b), our method achieves the best performance by balancing training stability and progressive preference integration.
Figure 2. Ablation study results of metric HR@50 on all datasets.
5.4. Hyper-Parameter Analysis
This subsection investigates the impact of several key hyper-parameters on model performance, including the number of user groups $g$, the number of local training epochs per client $H$, and the client sampling ratio per round $\beta$. For the number of user groups $\bm{g}$, experimental results reveal that model performance is sensitive to the number of user groups, with the best performance varying across datasets. An appropriate group number enables effective user partitioning and better captures preference diversity. For the number of local training epochs $\bm{H}$, setting local epochs to 5 generally yields the best performance. Fewer epochs may lead to underfitting, while excessive training can cause overfitting due to data heterogeneity. For the client sampling ratio $\bm{\beta}$, model performance remains generally comparable across different sampling ratios. Lower ratios may lead to slower convergence, whereas higher ratios help accelerate training and improve resource efficiency. Complete results and analysis on all datasets can be found in Appendix.
Figure 3. Visualization of client grouping dynamics during training on dataset Tools_and_Home. Each color indicates a distinct group.
5.5. Visualization Analysis
To better understand how our model organizes users into groups, we visualize the grouping dynamics during training on dataset Tools_and_Home. We randomly sample 100 users and track their group assignments across training rounds, with different colors indicating different groups. As shown in Figure 3, user groupings fluctuate frequently in early stages but gradually stabilize as training progresses. This indicates that the model captures latent user affinities, clustering users with similar characteristics into consistent groups. These results align with real-world scenarios, where different user cohorts exhibit distinct multimodal preferences, and support our motivation of leveraging group-level preferences to enhance the role of multimodal signals in recommendation.
5.6. Privacy Protection-Enhanced GFMFR
To further enhance the privacy protection of the proposed model, we integrate local differential privacy (LDP) by injecting Laplacian noise into client-side model parameters. Specifically, we experiment with five levels of noise intensity, setting the noise scale to $\delta=\{1,2,3,4,5\}$, approximating LDP settings under unit sensitivity, with injected noise levels corresponding to decreasing the privacy budget values in the range $[1,0.2]$. This setup allows us to analyze the trade-off between privacy and performance under different LDP configurations. As shown in Table 2, increasing the noise intensity improves privacy protection but leads to a gradual decline in model accuracy. In practical deployments, a moderate noise level can serve as a balanced trade-off, providing sufficient privacy without significantly sacrificing recommendation quality.
Table 2. Performance of integrating LDP into our proposed GFMFR under different noise scales. “H@50” and “N@50” are abbreviations for HR@50 and NDCG@50.
Dataset $\delta=0$ $\delta=1$ $\delta=2$ $\delta=3$ $\delta=4$ $\delta=5$
H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50
Beauty 0.2024 0.0730 0.1981 0.0633 0.1798 0.0591 0.1726 0.0587 0.1631 0.0496 0.1566 0.0443
Tools_and_Home 0.2269 0.0612 0.2025 0.0538 0.1708 0.0442 0.1660 0.0394 0.1550 0.0371 0.1529 0.0334
Toys_and_Games 0.2979 0.2613 0.2897 0.2471 0.2861 0.2402 0.2841 0.2393 0.2814 0.2385 0.2775 0.2365
Digital_Music 0.2943 0.2820 0.2786 0.2594 0.2668 0.2489 0.2650 0.2431 0.2603 0.2394 0.2551 0.2263
Office_Products 0.0614 0.0198 0.0119 0.0031 0.0109 0.0027 0.0105 0.0026 0.0100 0.0025 0.0095 0.0023
6. Conclusion
To address the limitations of existing federated recommender systems in handling rich multimodal content, we propose GFMFR, a flexible and efficient framework that systematically integrates multimodal information into the FR pipeline. By offloading the heavy multimodal representation learning to the server side and equipping clients with compact, distilled representations, GFMFR effectively alleviates client-side computation and communication burdens. The introduction of group-aware multimodal aggregation allows for adaptive knowledge sharing among users with similar preferences, while the preference-guided distillation ensures that multimodal signals are aligned with personalized recommendation objectives. Extensive empirical studies demonstrate the superiority of GFMFR in both performance and system efficiency across multiple datasets.